9. For the following set of scores, find the value of each expresion: b. (EX)2 c. Σ(X-1) d. Σ(X-1)2
9. For the following set of scores, find the value of each expresion: a.Σ b. (EX)2 c. Σ(X-1) d. Σ(x-1)?
Answer #1

Using R code we can easily solve this.

Input

x <- c(3,2,5,1,3)
sum(x^2)
(sum(x))^2
sum(x-1)
sum((x-1)^2)

Output

> x <- c(3,2,5,1,3)
> sum(x^2)
[1] 48
> (sum(x))^2
[1] 196
> sum(x-1)
[1] 9
> sum((x-1)^2)
[1] 25

\sum X^2= 48

(\sum X)^2= 196

\sum (X-1)= 9

\sum (X-1)^2= 25